OpenStudy (anonymous):

For what values of l and w will the perimeter of R be the least? Give a geometric explanation. Be sure to include a graph with relevant points labeled.

OpenStudy (anonymous):

Here is what I have so far:(attachment)

OpenStudy (anonymous):

for the full question: This time rectangle R has varying length l and width w but with a constant area of 4 square feet. a) Express the perimeter P as a function of length l. What type of function is P? What is the domain of P? b) Describe the asymptotic behavior of P. What can you say about rectangle R because of this behavior? Could you have made a similar statement about R back in Task 1? c) For what values of l and w will the perimeter of R be the least? Give a geometric explanation. Be sure to include a graph with relevant points labeled.

OpenStudy (amistre64):

well, you have P = 2(l+w), using some calculus we can determine that the rate of change of the perimeter is P' = 2l' + 2w' since you have w determine in terms of l. then w' = -4/l^2, and l'=1 P' = 2 + 2(4)/l^2, this is smallest when P'=0 0 = 2 + 2(4)/l^2 0 = 1 + 4/l^2 but this is only 0 when l^2 = -2 which isnt possible in real values.

OpenStudy (amistre64):

but thats just my attempt at it :)

OpenStudy (anonymous):

I think you may have forgotten the negative on -4/l^2 when you were plugging it into the equation, if so wouldn't it be l^2=2

OpenStudy (amistre64):

... you are correct lol

OpenStudy (amistre64):

it looks then like a square gives the smallest perimeter for the greatest area.

OpenStudy (anonymous):

so technically 2 is my answer for the values of l and w?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

ok much obliged

OpenStudy (amistre64):

good luck